OpenStudy (anonymous):

write the expression as a single, natural logarithm 1. 3ln 3 + 3ln c 2. simplify lne^3 3. ln2 + lnx = 5 please help

OpenStudy (anonymous):

gotta remember my log rules

OpenStudy (anonymous):

http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

@twodoggs I'm just extremely horrible at math... if i see a problem i can barely comprehend.. so those instructions don't even make sense to me ):

OpenStudy (anonymous):

3ln 3 + 3ln c will become \[\ln3^3+\ln c^3\]

OpenStudy (anonymous):

by the rules I posted ln simply means \[\log_{e} \]

OpenStudy (anonymous):

are you with me?

OpenStudy (anonymous):

\[3^3=27\]

OpenStudy (anonymous):

so we get \[\ln27+\ln c^3 \]

OpenStudy (anonymous):

which by the second rule logb(mn) = logb(m) + logb(n) becomes \[\ln (27c^3)\]

OpenStudy (anonymous):

that should help you with the second 2

OpenStudy (anonymous):

did any of that make sense?